Drake’s Market just announced its next annual dividend will be $1.50 a share. It expects the dividends to grow by 1.8 percent annually forever. How much will one share of this stock be worth five years from now if the required return is 15.5 percent?

Group of answer choices

A. $11.97

B. $11.76

C. $14.14

D. $12.19

E. $13.79
Answer #1

Current price=D1/(Required return-Growth rate)

=1.5/(0.155-0.018)

=10.9489051(Approx).

P5=Current price*(1+Growth rate)^5

10.9489051*(1.018)^5

=$11.97(Approx).
Answer #2

P0= 1.5(1.018)^5/ 0.155-0.018 =11.97

answered by: Ayyas